HUBBLE SPACE TELESCOPE DAILY REPORT #2091 PERIOD COVERED: 0000Z (UTC) 03/30/98 - 0000Z (UTC) 03/31/98 Daily Status Report as of 090/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt 2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/CCD 7803 (Daily Darks to Update Acquisition Bad Pixel Table Part III) The Space Telescope Imaging Spectrograph (CCD) was used to make daily dark calibration updates. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were completed as planned, and no anomalies were reported. 1.4 Completed Three Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed FOC/96 7347 (Direct Imaging of Betelgeuse) The Faint Object Camera (f/96) was used to make observations of the super giant star Betelgeuse (a.k.a., BD+75D325 or Alpha Orionis). Our UV images of Alpha Orionis obtained with the FOC in 1995 {Cycle 5} are the first direct images of the surface of a star other than the Sun. The observations were completed as planned, and no anomalies were reported. 1.6 Completed Eight Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.7 Completed STIS/CCD 7911 (STIS Galactic Plane Pure Parallel Program) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude survey observations in parallel with the prime science observations. Short exposures with the 50CCD imaging mode will provide the wavelength zero-point calibration. These data will be placed immediately into the Hubble Data Archive. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Three Sets of WF/PC-2 7625 (WF/PC-2 Cycle 7 Earth Flats) The WF/PC-2 was used to make a series of Earth flat calibration images. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The observations were completed as planned, and no anomalies were reported. 1.9 Completed NIC/1/2 7690 (NICMOS Pointed Super Flats) The Near Infrared Camera (Cameras One and Two) was used to make super flat field calibration observations. The observations will be used to build high signal/noise ratio NICMOS flat fields for the different filters. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Two Sets of NIC/2 7330 (The Fueling of Active Nucleii: A NICMOS Snapshot Survey) The Near Infrared Camera (Camera Two) was used to probe the dark matter distribution of NGC-4117 and NGC-6951 far from the center by observing intervening absorption from the Mg II doublet (lambda 2800) in the spectrum of one background QSO located in a geometrically optimal position. HST observations are required because this feature is in a wavelength range not accessible from the ground, and the QSO magnitudes and expected strength of absorption make IUE observations unfeasible. The proposal completed as expected with no anomalies. 1.11 Completed NIC/1/2/3 7899 (An IR Search For faint Companions to Pleiades Very Low-Mass Stars and Brown Dwarfs) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of CFHT-PL-20. This is part of a infrared campaign to search for faint companions to Pleiades very low-mass stars and brown dwarfs. The Pleiades has been the best site for brown dwarf {BD} searches, with half of the presently "confirmed" BDs being Pleiades members. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed Five Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.13 Completed WF/PC-2 7619 (WF/PC-2 Cycle 7 Decontamination) The WF/PC-2 was used to conduct a Cycle 7 decontamination procedure. Ultraviolet-blocking contaminants were removed, and hot pixels cured, by warming the CCDs to +20C for 6 hours. A variety of internal images were taken before and after the decontamination in order to verify its success and to monitor the instrument health. The procedures were conducted as scheduled, and no problems were reported. 1.14 Completed WF/PC-2 6856 (Early Evolution of Planetary mz-3 Nebulae) The WF/PC-2 was used to make observations of the MZ-3 class planetary nebula called IRC+10216 in order to study its early evolution. It is believed that the physical phenomena which shape the PNe occur early in their formation history. The observations were completed as planned, and no problems were reported. 1.15 Completed WF/PC-2 6619 (High Resolution Images of QSO Lyman Alpha Absorbing Galaxies) The WF/PC-2 was used to make several images of galaxy field (Field 1219) in order to obtain high resolution observations of the Lyman Alpha absorbing galaxies within this field-of-view. The purpose is to directly identify the objects responsible for low and intermediate redshift QSO Lyman-Alpha systems. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed STIS/MA1 7434 (The Supernova Intensive Study) The Space Telescope Imaging Spectrograph (MA1) was used to continue the long-term project on supernovae by observing SN1998S. HST can make unique contributions to understanding supernovae, as has been demonstrated through this program in past years, and now STIS and NICMOS offer new opportunities for deepening this investigation. The aim of the proposal is to explore the UV emission from supernovae, exploit the spatial resolution of HST and press the late-time observations of supernovae into uncharted territory. The observations were completed as planned, and no anomalies were reported. 1.17 Completed STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled, and no problems were reported. 1.18 Completed NIC/2 7227 (A Search for Low Mass/Sub-Luminous Companions to M-Stars) The Near Infrared Camera (Camera Two) was used to make a search for low mass/sub-luminous companions to low mass M class stars. Knowledge of stellar and sub-stellar masses and luminosities at and below the ~0.08 M{sun} H-burning limit is of fundamental importance. A dark calibration image was also taken. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 6 Successful : 6 2.2 FHST Updates: Scheduled : 17 Successful : 17 2.3 Operations Notes: The SSR EDAC error counter was cleared at 089/1653Z and at 090/0948Z per ROP SR-1. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH